The youngest of 4 children has siblings who are 3, 5, and 8 years older than she is. If the average (arithmetic mean) age of the 4 siblings is 32, what is the age of the youngest sibling?
Answer by VFBundy(438)

Age of youngest sibling = x
Age of 3rd oldest sibling = x + 3
Age of 2nd oldest sibling = x + 5
Age of oldest sibling = x + 8

x + (x + 3) + (x + 5) + (x + 8) = 4(32)

4x + 16 = 128

4x = 112

x = 28

Age of youngest sibling = x = 28